OpenStudy (anonymous):

how to find the derivative of y=3x-tanx

OpenStudy (anonymous):

3x has a very simple derivative. Tangent is a little more complicated, but it can be done with trig identities without too much trouble: \[\tan x=\frac{\sin x}{\cos x}\]

OpenStudy (anonymous):

Then use the quotient rule, and don't forget the - sign attached to the tangent!